OpenStudy (anonymous):

Find the coordinates of the center of the following circle. (x - 3)² + (y - 7)² = 49

OpenStudy (anonymous):

A circle with center \((x_0,y_0)\) and radius \(r\) has the equation \((x-x_0)^2+(y-y_0)^2=r^2\). So, taking that into consideration, what is the center of your circle?

OpenStudy (da_scienceman):

use d hint given and compare ur equationn with that of the standard equation for a circle with center x0,y0 and radius , r

OpenStudy (anonymous):

Select interior, exterior, or on the circle (x - 5) 2 + (y + 3) 2 = 25 for the following point

OpenStudy (anonymous):

(0, - 3)